OpenStudy (foxybear154):

Can someone please help me understand how to do this? Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (mathmale):

1. Find the slope of the line thru the given points. 2. Insert this slope into the point-slope form of the equation of a straight line:\[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

3. Chose either one of the given points. Take its coordinates and plug them into \[(x _{1},y _{1}).\]

OpenStudy (mathmale):

If done correctly, your result will be the desired answer!

OpenStudy (wolf1728):

Slope = m = (Y2 -Y1) / (X2 -X1) Slope = m = (-4 -0) / (0 --4) Slope = m = -4 / 4 Slope = m = -1 That's the slope. To generate the equation use y = mx + b y = -x + b Use one of the given points (0, 0) and put those numbers into the formula.